A 50kg mass is suspended from one end of a wire of length 4m and diameter of 3mm whose other end is fixed. What will be the elongation of the wire? Take the young modulus of the material to be 7 × 10^10 Nm^-2
1
Expert's answer
2021-09-12T19:00:37-0400

According to the Hooke's law, the elongation is given as follows:


"E\dfrac{\Delta l}{l} = \sigma\\
\Delta l =\dfrac{ \sigma}{lE}"

where "l" "=4m" is the length of the wire, "E = 7\times 10^{10}Nm^{-2}" is the young modulus of the material, and "\sigma" is the stress in the wire. By definition, this stress is given as follows:


"\sigma = \dfrac{F}{S}"

where "F = mg" is the weight of the mass "(m=50kg, g = 9.8N\/kg)", and "S = \pi d^2/4" is the cross-sectional area of the wire ("d = 3mm = 3\times 10^{-3}m" is the diameter of the wire).

Putting it all together, obtain:


"\Delta l = \dfrac{4mgl}{\pi d^2E}\\
\Delta l = \dfrac{4\cdot 50kg\cdot9.8N/kg\cdot 4m}{\pi \cdot (3\times 10^{-3}m)^2\cdot 7\times 10^{10}N/m^2} \approx 3.96\times 10^{5}m"

Answer. "3.96\times 10^{5}m".